# url]https://person.dibris.unige.it/ricca-filippo/

### v1

@@ 6.2.4 Open Challenges of AIaTA
As such, our participants emphasized developing self-check mechanisms and best practices for using AI in TA. LLMs often attempt to create solutions rather than acknowledging uncertainty. LLMs have a user-pleasing tendency, finding workarounds instead of reporting errors (obviating the oracle problem). This can lead to excessively positive results, while testers typically adopt a more cautious, critical, and pessimistic perspective. Our interviews revealed that quality is not just about coverage; relevance and better assertions are crucial to avoid potential risks in production.
@@ 7.3 Secondary Studies on AI in Test Automation
Our work falls within the scope of these secondary studies. Unlike previous works, we focused on the grey literature to gain insights into practitioners’ perceptions of AI/ML adoption in TA. Moreover, we validated and extended our taxonomies with interviews with industrial professionals.
@@ 8 Conclusions and Future Work
To fill this gap, in this paper, we present a multi-year study of the grey literature concerning AI solutions for test automation. We manually analyzed several thousands of documents from which we retrieved many problems about different aspects of the automated testing process. Moreover, our taxonomy includes the solutions that are used to mitigate such problems and the list of the most popular tools available. Our taxonomies were validated by five interviews with industrial practitioners, who also provided further insights about the usage of AI in TA.
@@ 9 Acknowledgments
We extend our deepest gratitude to the professionals who generously took the time to participate in our interviews.

### v2

@@ 8.2 Observations from the Interviews
As such, our participants emphasized developing self-check mechanisms and best practices for using AI in TA. LLMs often attempt to create solutions rather than acknowledging uncertainty. LLMs have a user-pleasing tendency, finding workarounds instead of reporting errors (obviating the oracle problem). This can lead to excessively positive results, while testers typically adopt a more cautious, critical, and pessimistic perspective. Our interviews revealed that quality is not just about coverage; relevance and better assertions are crucial to avoid potential risks in production. In practice, the testing team creates test suites exhibiting “a reasonable coverage for their needs and for their criticality”.
@@ 9.3 Secondary Studies on AI in Test Automation
Our work falls within the scope of these secondary studies. Unlike previous works, we focused on the grey literature to gain insights into practitioners’ perceptions of AI/ML adoption in TA. Moreover, we corroborated our taxonomies with interviews with industrial professionals.
@@ 10 Conclusions and Future Work
To fill this gap, in this paper, we present a multi-year study of the grey literature concerning AI solutions for test automation. We manually analyzed several thousands of documents from which we retrieved many problems about different aspects of the automated testing process. Moreover, our taxonomy includes the solutions that are used to mitigate such problems and the list of the most popular tools available. The insights provided by our taxonomies were corroborated by five interviews with industrial practitioners, who also provided further insights about the usage of AI in TA.
@@ 11 Acknowledgments
This research was funded by the Bavarian Ministry of Economic Affairs, Regional Development and Energy. We extend our deepest gratitude to the professionals who generously took the time to participate in our interviews.